Find equation tangent to parabola y^2=4x whose intercept on y axis is 2

Grade:11

1 Answers

Arun
6 years ago
Dear Aditya
In y² = 4x, a = 1
 
Tangent to parabola
y = mx + (a/m)
 
But a/m = 2
a = 1
m = ½
Now
Tangent equation will be
Y = (1/2)x + 2
2y = x +4
 
Regards
Arun (askIITians forum expert)